### v1

@@ 5.4 Kinematics
Despite the unknown DIB carriers, many works have attempted to use DIB as a tracer to study the kinematics of the ISM since the DIB carriers are proven to be interstellar. Zasowski et al. (2015) provided what 3D distribution the carriers of the DIB at 1.5273 $\mu$m have, and were first to elaborately depict how these DIB carriers move with the Milky Way by the mean Galactic velocity curve. Following this work, Zhao et al. (2021b) and Gaia Collaboration et al. (2023b) displayed the kinematics of the DIB $\lambda 8621$ in the same way, while the average galactocentric distance of the DIB $\lambda 8621$ estimated by the velocity curve is about 7.5 kpc, smaller than 9 kpc derived from Zasowski et al. (2015) due to the differences of the distances of the observed stars. In addition, a comparison of the LSR velocity of the DIB $\lambda 8621$ with that of CO gas (Dame et al. 2001) shows they own the same kinematics pattern, which was quantified by the linear correlation of a slope of 0.95 using the algorithm called “peak-finding” (Saydjari et al. 2023). Inspired by these works, we investigate the LSR velocity of the DIBs $\lambda 5780$ and $\lambda 6614$, $V_{\rm DIB}^{\rm LSR}$, as a function of the Galactic longitude $\ell$, as presented in Figure 13. The comparison of $V_{\rm DIB}^{\rm LSR}$ with the LSR velocity of the CO gas and the other tracers of the ISM will be explored in the forthcoming work.
@@
where $c$ is the speed of light, $\lambda_{\rm obs}$ is the observed central wavelength of DIB, $\lambda_{0}$ is the final central wavelength of DIB fitted by the statistical method in Section 5.3, and $\hat{r}$ is the unit direction vector from the Sun to the source. The Galactic velocity curves with diverse heliocentric distances $d$ are computed using the Model A5 of Reid et al. (2019). It can be seen from Figure 13 that the fluctuation of $V_{\rm DIB}^{\rm LSR}$ for the whole selected samples reflects the motion of the DIB carrier with the Galactic rotation. The medians of $V_{\rm DIB}^{\rm LSR}$ in each bin ($\Delta\ell=10^{\circ}$) are both consistent with the velocity curve with the heliocentric distance $d=0.5\ \rm kpc$ at the $30^{\circ}<\ell<180^{\circ}$ and the $d=1.5\ \rm kpc$ at the $\ell>180^{\circ}$. The results agree with the distance of the DIB $\lambda 8621$ estimated Gaia Collaboration et al. (2023b), but may be related to the fact that most of the background stars are located within 3 kpc.
@@ 6 Data and Code Availability
The final catalog of the DIB measurements of the 7,681 common HQ sources, and the catalogs of the HQ and the MQ samples for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are publicly available at https://nadc.china-vo.org/res/r101404/ in FITS tables. The codes used to perform the pipeline of measuring DIB are also provided via https://github.com/iScottMark/LAMOST_DIB.
@@ 7 Summary and Conclusions
We successfully extract 2,188,240 ISM residual spectra from the late-type stellar spectra in the LAMOST LRS using the method of subtracting the stellar templates. By fitting the ISM residual spectrum with the Gaussian profile, the largest sample to date of reliable measurements of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in the northern sky is obtained. The selected 176,831, 13,473 and 110,152 HQ samples of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively, corresponding to 142,074, 11,480, and 85,301 unique sources, not only allow for a more comprehensive review of DIB correlations from a statistical perspective, but also provide the first characterization of the sky distribution and kinematic properties of these three DIBs. Besides, we also release a total of 27,598, 42,735 and 21,635 MQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively, corresponding to 23,920, 36,825 and 18,145 unique stars. These marginal or null measurements allow for further studies such as 3D mapping or additional correlations. Our primary conclusions are as follows:
(i) By virtue of the largest sample size of the HQ measurements of the DIBs $\lambda 5780$, $\lambda 6614$, it is the first time to portray their Galactic distribution at a resolution of 0.92/pixel.
(ii) In our statistical analysis with a sample size on the order of tens of thousands, we provide more consistent and dependable EW of DIB per unit $\rm E(B-V)$. For the DIB $\lambda 5780$, the value is determined to be 0.565 Å$\rm\AA\,mag^{-1}$, which is larger than previous works, while for the DIBs $\lambda 5797$ and $\lambda 6614$, the values are 0.176 and 0.256 Å$\rm\AA\,mag^{-1}$, respectively, which agree with preceding findings.
(iii) Without assuming the co-motion of the DIB carriers and the ISM, we use a kinematic fitting method to get the precise measurements of the central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, which are $5780.48\pm 0.01$, $5796.94\pm 0.02$ and $6613.64\pm 0.01$, respectively.
@@ Acknowledgements.
This work is supported by National Key R&D Program of China (Grant NO. 2019YFA0405102) and National Natural Science Foundation of China (grant Nos. 12261141689, 11973060 and 12090044). Guoshoujing Telescope (the Large Sky Area Multi-Object Fiber Spectroscopic Telescope, LAMOST) is a National Major Scientific Project built by the Chinese Academy of Sciences. Funding for the Project has been provided by the National Development and Reform Commission. LAMOST is operated and managed by the National Astronomical Observatories, Chinese Academy of Sciences. He Zhao acknowledges the National Natural Science Foundation of China (grant No. 12203099), the China Postdoctoral Science Foundation (No. 2022M723373), and the Jiangsu Funding Program for Excellent Postdoctoral Talent.
@@ Appendix A Instrumental Broadening in LAMOST LRS
Compared with the high-resolution spectrographs, the low-resolution spectrographs have a more significant instrumental broadening effect. For the LAMOST LRS, the instrumental broadening effect is mainly caused by the slit width, $\rm\Delta\lambda$. Although Cui et al. (2012) and Luo et al. (2012) have reported the design of LAMOST including $\rm\Delta\lambda$, it was an experimental parameter at that time. The actual resolution of LAMOST LRS is about 1800 (private communication, Yonghui Hou, 2023) at 5500 Å in red end (from 3300 Å to 5900 Å) and at 7500 Å in blue end (from 5700 Å to 9000 Å). The resolution R is defined as $\rm\lambda/\Delta\lambda$, where $\rm\lambda$ is the corresponding central wavelength. Hence, $\rm\Delta\lambda$, i.e. the instrumental FWHM at 5500 Å and 7500 Å are roughly 3.06 Å and 4.17 Å.
@@
We refer to the canonical FWHM of DIB (see the column $\rm FWHM_{ref}$ in Table 6) from Vogrinčič et al. (2023) as our benchmark to estimate the practical Gaussian width of DIB against the LAMOST LRS. The broadened Gaussian width of DIB can be approximated by Equation 8.

### v2

@@ 5.4 Kinematics
Despite the unknown DIB carriers, many works have attempted to use DIB as a tracer to study the kinematics of the ISM since the DIB carriers are proven to be interstellar. Zasowski et al. (2015) provided what 3D distribution the carriers of the DIB at 1.5273 $\mu$m have, and were the first to elaborately depict how these DIB carriers move with the Milky Way by the mean Galactic velocity curve. Following this work, Zhao et al. (2021b) and Gaia Collaboration et al. (2023b) displayed the kinematics of the DIB $\lambda 8621$ in the same way, while the average galactocentric distance of the DIB $\lambda 8621$ estimated by the velocity curve is about 7.5 kpc, smaller than 9 kpc derived from Zasowski et al. (2015) due to the differences of the distances of the observed stars. In addition, a comparison of the LSR velocity of the DIB $\lambda 8621$ with that of CO gas (Dame et al. 2001) shows they own the same kinematics pattern, which was quantified by the linear correlation of a slope of 0.95 using the algorithm called “peak-finding” (Saydjari et al. 2023). Inspired by these works, we investigate the LSR velocity of the DIBs $\lambda 5780$ and $\lambda 6614$, $V_{\rm DIB}^{\rm LSR}$, as a function of the Galactic longitude $\ell$, as presented in Fig. 13. The comparison of $V_{\rm DIB}^{\rm LSR}$ with the LSR velocity of the CO gas and the other tracers of the ISM will be explored in the forthcoming work.
@@
where $c$ is the speed of light, $\lambda_{\rm obs}$ is the observed central wavelength of DIB, $\lambda_{0}$ is the final central wavelength of DIB fitted by the statistical method in Sect. 5.3, and $\hat{r}$ is the unit direction vector from the Sun to the source. The Galactic velocity curves with diverse heliocentric distances $d$ are computed using the Model A5 of Reid et al. (2019). It can be seen from Fig. 13 that the fluctuation of $V_{\rm DIB}^{\rm LSR}$ for the whole selected samples reflects the motion of the DIB carrier with the Galactic rotation. The medians of $V_{\rm DIB}^{\rm LSR}$ in each bin ($\Delta\ell=10^{\circ}$) are both consistent with the velocity curve with the heliocentric distance $d=0.5\ \rm kpc$ at the $30^{\circ}<\ell<180^{\circ}$ and the $d=1.5\ \rm kpc$ at the $\ell>180^{\circ}$. The results agree with the distance of the DIB $\lambda 8621$ estimated Gaia Collaboration et al. (2023b), but may be related to the fact that most of the background stars are located within 3 kpc.
@@ 6 Data and code availability
The final catalog of the DIB measurements of the 7,681 common HQ samples, and the catalogs of the HQ and the MQ samples for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are publicly available at https://nadc.china-vo.org/res/r101404/ in FITS tables. The codes used to perform the pipeline of measuring DIB are also provided via https://github.com/iScottMark/LAMOST_DIB.
@@ 7 Summary and conclusions
We successfully extracted 2,188,240 ISM residual spectra from the late-type stellar spectra in the LAMOST LRS using the method of subtracting the stellar templates. By fitting the ISM residual spectrum with the Gaussian profile, we obtained the largest sample of reliable measurements of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in the northern sky to date. The selected 176,831, 13,473, and 110,152 HQ samples of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively, corresponding to 142,074, 11,480, and 85,301 unique sources, not only allow for a more comprehensive review of DIB correlations from a statistical perspective, but also provide the first characterization of the sky distribution and kinematic properties of these three DIBs. In addition, we release a total of 27,598, 42,735, and 21,635 MQ measurements of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively, corresponding to 23,920, 36,825, and 18,145 unique stars. These marginal or null measurements allow further studies, such as 3D mapping or the investigation of additional correlations. Our primary conclusions are as follows:
(i) We release to the community the largest sample of HQ measurements of the DIBs $\lambda 5780$ and $\lambda 6614$ to date, and by virtue of these data, we also present their Galactic distribution for the first time, at a resolution of 0.92/pixel.
(ii) In our statistical analysis with a sample size on the order of tens of thousands, we provide more consistent and robust EWs of DIBs per unit $\rm E(B-V)$. For the DIB $\lambda 5780$, we determine a value of 0.565 Å$\rm\AA\,mag^{-1}$, which is larger than those found in previous works, while for the DIBs $\lambda 5797$ and $\lambda 6614$, we obtain values of 0.176 and 0.256 Å$\rm\AA\,mag^{-1}$, respectively, which agree with previous findings.
(iii) Without assuming the co-motion of the DIB carriers and the ISM, we use a kinematic fitting method to obtain precise measurements of the central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, which are $5780.48\pm 0.01$, $5796.94\pm 0.02$, and $6613.64\pm 0.01$, respectively.
@@ Acknowledgements.
This work is supported by National Key R&D Program of China (Grant NO. 2019YFA0405102) and National Natural Science Foundation of China (grant Nos. 12261141689, 11973060, and 12090044). Guoshoujing Telescope (the Large Sky Area Multi-Object Fiber Spectroscopic Telescope, LAMOST) is a National Major Scientific Project built by the Chinese Academy of Sciences. Funding for the Project has been provided by the National Development and Reform Commission. LAMOST is operated and managed by the National Astronomical Observatories, Chinese Academy of Sciences. He Zhao acknowledges the National Natural Science Foundation of China (grant No. 12203099), the China Postdoctoral Science Foundation (No. 2022M723373), and the Jiangsu Funding Program for Excellent Postdoctoral Talent.
@@ Appendix A Instrumental broadening in LAMOST LRS
Compared with the high-resolution spectrographs, the low-resolution spectrographs have a more significant instrumental broadening effect. For the LAMOST LRS, the instrumental broadening effect is mainly caused by the slit width, $\rm\Delta\lambda$. Although Cui et al. (2012) and Luo et al. (2012) have reported the design of LAMOST including $\rm\Delta\lambda$, it was an experimental parameter at that time. The actual resolution of LAMOST LRS is about 1800 (private communication, Yonghui Hou, 2023) at 5500 Å in red end (from 3300 Å to 5900 Å) and at 7500 Å in blue end (from 5700 Å to 9000 Å). The resolution R is defined as $\rm\lambda/\Delta\lambda$, where $\rm\lambda$ is the corresponding central wavelength. Hence, $\rm\Delta\lambda$, namely the instrumental FWHM, at 5500 Å and 7500 Å are roughly 3.06 Å and 4.17 Å.
We refer to the canonical FWHM of DIB (see the column $\rm FWHM_{ref}$ in Table 6) from Vogrinčič et al. (2023) as our benchmark to estimate the practical Gaussian width of DIB against the LAMOST LRS. The broadened Gaussian width of DIB can be approximated by Eq. (8).